OpenStudy (anonymous):

4^2x+3 =1 and 2^2x+5-8x

OpenStudy (anonymous):

\[\LARGE 4^2x+3=1\] or \[\LARGE 4^{2x+3}=1\] and what do we have to do with the second expression ? :)

OpenStudy (anonymous):

the second one and we just have to simplify, but I have no idea how to

OpenStudy (anonymous):

if you simplyfy this is the answer

OpenStudy (anonymous):

5-4x

OpenStudy (anonymous):

for the # 2

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

ok... regarding to the equation \[\LARGE 4^{2x+3}=1\] we take logs to both sides... \[\LARGE \log(4)^{2x+3}=\log (1)\] \[\LARGE (2x+3)\cdot \log(4)=0\] now we have two cases \log(4)=0 ---which is not true. or to solve 2x+3=0 2x=-3 x=-3/2 Which is true... because: \[\LARGE 4^{2x+3}=1\] \[\LARGE 4^{2\cdot (-3/2)+3}=1\] \[\LARGE 4^0=1\] 1=1 so \[\LARGE x=-\frac32 \]

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

@reticens your second expression is it like: \[\LARGE 2^2x+5-8x\] or \[\LARGE 2^{2x+5-8x}\] ?

OpenStudy (anonymous):

the second one except it is =8x and the 8x is equal for all of the equation

OpenStudy (anonymous):

\[\LARGE 2^{2x+5}=8x\] this?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

then forget what jorge said, he meant another way :) ... \[\LARGE 2^{2x+5}=8x\] it's complicated ...let me think :(

OpenStudy (anonymous):

alright thank you :)

OpenStudy (anonymous):

and yes i know :(

OpenStudy (anonymous):

well I have to confess that I'm not capable to solve it. looks like every value of X we want to take... we'll never get the equation right... :(

OpenStudy (anonymous):

its okay! I'll just make a guess haha

OpenStudy (anonymous):

I'm affraid you can't.... anyway Good Luck, or hold on. Call someone smarter, maybe they can help you out.

OpenStudy (anonymous):

alright thank you!